(D3 − 3D2 − 6D + 8)y = xe−3x


1
Expert's answer
2021-04-20T16:37:03-0400

Given,

"(D^3 \u2212 3D^2 \u2212 6D + 8)y = xe^{\u22123x}"


Its auxilary equation is-


"m^3-3m^2-6m+8=0"


"m^3-m^2-2m^2+2m-8m+8=0\\

m^2(m-1)-2m(m-1)-8(m-1)=0\\

(m-1)(m^2-2m-8)=0"


The roots of the equation is-

"m=1,4,-2"


Complimentary function is-

"C.F.=c_1e^x+c_2e^{4x}+c_3e^{-2x}"


Particular integral

"P.I.=\dfrac{e^{-3x}x}{D^3-3D^2-6D+8}"


"=\dfrac{e^{-3x}x}{(D-3)^2-3(D-2)^2-6(D-1)+8}"


"=\dfrac{e^{-3x}x}{D^3-12D^2+33D-25}"


"=\dfrac{e^{-3x}x}{-25(1-(\dfrac{D^3-12D^2+33D)}{25})}"


"=\dfrac{e^{-3x}}{-25}(1-\dfrac{D^3-12D^2+33D}{25})x"


"=\dfrac{-e^{-3x}}{25}(x-\dfrac{33}{25})"


"=\dfrac{-e^{-3x}}{25}(x-\dfrac{33}{25})"


Complete solution is-


"y=c_1e^x+c_2e^{4x}+c_3e^{-2x}-\dfrac{e^{-3x}}{25}(x-\dfrac{33}{25})"